Software and Computing

Software and Computing Phone Meeting

2011-02-16 12:00
2011-02-16 13:00
America/New York
Wednesday, 16 February 2011
1-189, EVO, at 17:00 (GMT), duration : 01:00
TimeTalkPresenter
12:00Sti/VMC update ( 00:20 ) 1 fileVictor Perevoztchikov (BNL)
12:20AOB ( 00:20 ) 0 filesAll (All)

Run 11 preparation/support meeting #12

2011-02-11 14:00
2011-02-11 15:00
America/New York
Friday, 11 February 2011
1-189, EVO, at 19:00 (GMT), duration : 01:00

Software and Computing Phone Meeting

2011-02-09 12:00
2011-02-09 13:00
America/New York
Wednesday, 9 February 2011
1-189, EVO, at 17:00 (GMT), duration : 01:00
TimeTalkPresenter
12:00Embedding overview, path to QM 2011 ( 00:15 ) 1 fileRenee Fatemi (UKY)
12:15Run 11 initial calibration update ( 00:20 ) 1 fileGrant Webb (UKY)
12:35Voltage scan and comisc ray study ( 00:10 ) 0 filesYuri Fisyak (BNL)
12:45VMC geometry update ( 00:15 ) 1 fileJason Webb (BNL)

Run 11 preparation/support meeting #11

2011-02-04 14:00
2011-02-04 15:00
America/New York
Friday, 4 February 2011
2-187, EVO, at 19:00 (GMT), duration : 01:00

Minutes 

Attendees: Dmitry A., Wayne B., Leve H., Jérôme L., Gene V.B., Jan B., Maxim N.

 

Online webserver:

  • Wayne installed: SSH key client, osiris
  • ITD backups nominally working
    • Warning message noted to ITD
    • Restore from backup (targeting machine with warning messages) by Wayne went fine
  • Synchronization with secondary new webserver done manually, not yet automatic
  • Tomcat restarted for Leve's scheduler statistics redeployment
    • Jerome's service idea should be copied from old server so that intentional downtimes don't trip the watcher

 

ShiftLog:

  • Detector checklist being posted to the ShiftLog regularly now. A suggestion from Jan to make it easier was to make a shortcut to the checklist document from the ShiftLeader's desktop (post-meeting Leve learned that each new checklist is given a new document name; finding the document isn't that difficult anyhow)

 

Online user support:

  • Enabling migration off sc.starp and onto linux pool
    • Dmitry notes that Corba libs are present on linux pool
    • Jan says Brian Page is in the midst of trying to move two components:
      • cdev reader (needs Corba libs, will try compiling)
      • DB writer (Gene will provide Brian with a working example code)
  • Two online disks (noted by Wayne in previous meetings): replaced
  • Some additional network issues this week
    • Initially fixed with a switch replacement
    • ...but addition failures occurred
    • Wayne will try to fix during access later today (post-meeting update: this went successfully; topic closed)
  • starutilities replacement: usage extension currently with "soft" limit
    • Matt could not install C-AD software in XP mode on new box
    • next possibility is dedicated XP machine (needs a software extension from ITD)
  • Other tasks (noted in previous meetings) in queue

 

Databases:

  • Shift sign-up un-subscriptions now exclude trainees [Dmitry]
  • Feedback to slow controls to implement monitoring of their automated services (are they running? is data flowing? for now, only Dmitry's dbPlots and an environment monitor)
  • TPC anode voltage migration was not running, but Dmitry will turn it on (we are aware of its shortcomings)
  • Dmitry noted RHIC clock seems to be variable even during fills (Gene will follow up)
  • Some discussion of when DB snapshots run; Jan seems OK with packaging at midnight for 4am usage for cloud jobs
  • Dmitry notes that there's an open discussion about recorded prescales being wrong. If this is corrected later in the DB, it shouldn't be a problem for offline users as long as they access from the DB and not from the DAQ files. Needs to be sorted out by trigger group first.

 

FastOffline:

  • Laser files were being skipped; fixed by Jerome
  • Automatic drift velocity calibrations started (and working)

 

QA:

  • OnlinePlots had 3 problems spurred by update of starnew to SL11a library
    • Macros used (Load.C) needed modified loading order of libraries (fixed)
    • Numerous histograms getting skipped, and others in unusual order in output PDFs (RT 2085)
      • Temporary solution: run in starpro (SL10k) patched for TOF/VPD updates
    • Code for inserting PDFs into database should use 'REPLACE INTO' instead of 'INSERT INTO', but we only have a binary of the current version
      • Dmitry had source code from an old version which Gene edited to work as desired now (update: codes committed to CVS)
  • Offline QA histogram filling needed to update code for TPC padrow determination
  • AutoCombining of offline histograms stopped working (Elizabeth investigating) (update: fixed for a changed function interface in QA daemon)
  • No references yet, but shift crew (Lanny) looking at data
    • Call for setting references will go out when physics readiness is achieved (many problems in current data)

 

Run 11 preparation meeting #10

2011-01-28 14:00
2011-01-28 15:00
America/New York
Friday, 28 January 2011
1-189, EVO, at 19:00 (GMT), duration : 01:00

Minutes 

Attendees: Dmitry A., Wayne B., Leve H., Jérôme L., Gene V.B.

 

Online webserver:

  • Went live on Monday
  • No user complaints reported
  • Jerome: identified and fixed several problems
  • Jerome: started tomcat watcher script (he noted that it already caught a number of trips)
  • Dmitry: though crontab had been copied, cron file entries had been missed
  • Wayne: to-do list:
    • Install SSH key client
    • Install osiris
    • Tomcat log file rotation
    • ITD backups (noted as priority)
    • Synchronization with secondary new webserver

ShiftLog:

  • 'thumbnailer' problem, Jerome is helping
  • MTD subsystem added by Leve
    • Used the opportunity to move the array of subsystems to a single place

 

Online user support:

  • sc.starp crashes (RT 2077)
    • Was repeatedly shutting down
    • Wayne: Box replaced (drive swapped), working fine...
    • ...but suspicion is UPS, battery replaced
    • Leave as is for now
    • Dmitry suggested providing Corba libs on online linux pool nodes to allow users an alternative to sc (Gene spoke with Jan B. about this post-meeting: Brian P. will try migrating his code)
  • Two online disks (noted by Wayne in previous meetings): still need replacement
  • QtRoot Event Display (RT 2075): Jerome notes it is low priority for now
  • HPSS access from online via data carousel: Jerome indicates work in progress for authentication (perhaps try replicating what was done with stargrid04)
  • starutilities replacement: extension through Jan. 30th.
    • trying to run C-AD software in XP mode on new box (Matt working with ITD today)
    • next possibility is dedicated XP machine
    • otherwise suggest software upgrade to C-AD (would require additional extension for procurement and delivery duration)
  • rts01 replacement machine (RT 2070): machine has arrived
  • Request for help restoring old FTPC event builders (RT 2066):
    • Wayne attempted to resurrect old builders but didn't succeed
    • Jeff has alternatives to move on
  • Conference room PC display configuration improved by Wayne, but may still be not optimal
  • FMS/Spin request for replacement machines (RT 2073 and 2074):
    • Laptop provided, and had to learn from experience to prevent power saver settings from shutting it down
    • Other machine not critical
  • Renewed request today for a folder on ShiftLeader's terminal for making files web-available (RT 2083)

 

Databases:

  • Migration macros started on onl01 by Dmitry
    • Machine died (from exhausted memory): cause is unknown, but perhaps one of the users
      • Table further investigation until it proves to be systematic
  • New DB node allocation for FatOffline(+nightly tests) awaits results of Lidia's tests
  • Dmitry noted that detector ID lists needed to be (and were) verified among Lidia, Tonko, and himself
    • Jerome noted that offline should rely on DB lookup by subsystem and not depend on a local list (he will follow up with Lidia)
  • Dmitry implemented flexible, graphical selection of timestamps for his online DB plots upon feedback from Alexei L.
    • Jerome tested on several OS-browser combinations

 

FastOffline:

  • Jerome notes that the chain needs continued attention (e.g. geometry, CA tracker), and Gene will take this

 

QA:

  • Some minor modifications made to OnlinePlots for VPD by Kefeng and Gene

 

Software and Computing Phone Meeting

2011-02-02 12:00
2011-02-02 13:00
America/New York
Wednesday, 2 February 2011
1-189, EVO, at 17:00 (GMT), duration : 01:00
TimeTalkPresenter
12:00FMS simulation production, status ( 00:15 ) 0 filesJason Webb (BNL)
12:15Production status & step forward ( 00:20 ) 0 filesJerome Lauret (BNL)
12:35AOB ( 00:20 ) 0 filesAll (All)

Run 11 preparation meeting #9

2011-01-21 14:00
2011-01-21 15:00
America/New York
Friday, 21 January 2011
1-189, EVO, at 19:00 (GMT), duration : 01:00

Minutes 

Attendees: Dmitry A., Wayne B., Leve H., Jérôme L., Gene V.B.

 

ShiftLog [Leve]:

  • Checked that links work both ways with RunLog
  • Re-deployed to fix an issue with a function used to tell if we're in a shift-taking period
    • Dmitry suggested to modify this next year to not use dates

 

Online infrastructure [Wayne]:

  • New webserver:
    • Database access now working fine
    • Watcher script needs turned on
    • cron jobs (and associated files) need to be copied
    • Announcement for a Monday swap:
      • 2pm (Wayne, Jerome, Leve, Dmitry should be available to handle whatever may arise - Dmitry already knows one configuration he needs to change during the swap)
      • NFS will be turned off on the old server, sync will be done, IPs swapped, and NFS on.
  • Two online disks (noted last week) still need replacement
  • Event Display
    • Not currently functional (we lost our maintainer; Jerome is trying to prop it back up for now)
    • Pieces in multiple locations for online display plus offline (images dumped for web display)
    • Not critical (old online event display still functional), but valuable for PR
  • Long-standing request (not critical) to be able to read from HPSS online (via data carousel)
    • Jerome looking for any road blocks (e.g. firewall)
  • Machine replacements:
    • emcsc done
    • starutilities and ftpcgas not yet done (scheduled for work with relevant parties early next week)
  • All Windows nodes rebooted with security patches
  • rts01 replacement request merged into a larger purchase request (RT 2070)
  • Request for help restoring old FTPC event builders (RT 2066)
    • Possible solution of other hardware if this is too much effort
  • FMS/Spin request for replacement machines (RT 2073 and 2074). Not critical, but latter is important for remote access (current machine is inside WAH and not on network)

 

Databases [Dmitry]:

  • After new nodes go into production pool, 2 old nodes will be targeted for FastOffline + nightly tests
    • Assessment of new node speed/power still ongoing
  • onldb1 and 2 were up for 1+ years, so rebooted on purpose to check state
    • Collector daemons did not automatically start, but otherwise OK
    • Dmitry suggested leaving it this way since reboot need is so rare and automatic can take a while to get caught up (has to look through and skip all runs from the year)
    • Gene suggested making it automatic in case Dmitry's not around to restart daemons (though shift leaders can potentially do it themselves). We'll go with this.
  • QA files (OnlinePlots + Jevp) stored for RunLog can now exceed 4 GB without needing manual rollover
  • TPC Anode Voltges:
    • For now, continue with methods of collection from last year
    • Dmitry tested a supposedly-optimized "bulk reading" methods and found them to be no better than channel-by-channel (actually worse as it appears to be the same + overhead: 13-15 seconds to collect all values versus less than 1 second); dead end.
    • Possible new methods of collection under development (Tonko & the TPC group)
    • Feedback to the TPC group that wider "dead band zones" could reduce traffic and help alleviate problems due to load

 

FastOffline [Jerome]:

  • Same as last week: many files coming into the system, but all are tests, pulsers, etc., so not being processed
  • Gene notes some possible tweaks to the chain may be needed for TPC calibration status (awaiting TPC feedback)
  • Dmitry need to verify with Lidia some exported info from from onldb1

 

QA [Gene]:

  • Offline QA documentation updated (Lanny & Gene)
  • Jerome notes the need to monitor and document and issues with QA Shift
  • Pre-combining of FastOffline files still needs tested and configured to be the default

 

Software and Computing Phone Meeting

2011-01-19 12:00
2011-01-19 13:00
America/New York
Wednesday, 19 January 2011
1-189, EVO, at 17:00 (GMT), duration : 01:00
TimeTalkPresenter
12:00Simulation production overview ( 00:20 ) 0 filesJason Webb
12:20TPC Alignement revisit ( 00:10 ) 0 filesYuri Fisyak (BNL)
12:30AOB ( 00:20 ) 0 filesAll (All)

TPC Voltage issue and discussion #3

2011-01-18 13:00
2011-01-18 14:00
America/New York
Tuesday, 18 January 2011
1-189, at 18:00 (GMT), duration : 01:00

Meetings are being set to discuss and converge as far as the TPC voltage issue is concerned.

This is the third of this kind.

 

TPC Voltage issue and discussion #2

2011-01-18 10:00
2011-01-18 11:00
America/New York
Tuesday, 18 January 2011
1-189, at 15:00 (GMT), duration : 01:00

This meeting is the second of this kind. Meetings are being set to discuss and converge as far as the TPC voltage issue is concerned.

Several action items and discussions were started since the first meeting (see especially this thread).

 

TPC Voltage issue and discussion #1

2011-01-14 16:30
2011-01-14 17:30
America/New York
Friday, 14 January 2011
1-189, at 21:30 (GMT), duration : 01:00

Meeting was set to discuss on was to converge with the TPC voltage issue. This is the first of a serie of meetings with the diverse experts.

Run 11 preparation meeting #8

2011-01-14 14:00
2011-01-14 15:00
America/New York
Friday, 14 January 2011
1-189, EVO, at 19:00 (GMT), duration : 01:00

Minutes 

Attendees: Wayne B., Leve H., Jérôme L., Gene V.B.

 

ShiftLog [Leve]:

  • New server:
    • Ready:
      • App copied to new server (Wayne)
        • Successfully passed stability tests (pounding on it)
      • All WARs deployed
      • Scripts ready to flush cache and re-copy WARs
    • To do:
      • Will re-check links to and from RunLog
      • Watcher script not on yet
      • Verify Offline QA Shift Report connections

 

Online infrastructure [Wayne]:

  • New web server:
    • Links on homepage checked
      • Some issues with DB access and privilege from current node name (dean2)
    • New jpgraph version installed in same location as the old one
    • Disk exports are not in place, so content isn't being updated currently
    • Will do another sync over the weekend or early next week, in addition to the final one when we swap servers
    • Ganglia to be installed
    • User accounts to be added
  • New web server backup:
    • Configuration kept the same as main new web server
    • ...but content isn't sync-ed
  • Password rotation completed
  • mysql-devel installed on online linux nodes (RT 2068)
  • Two online disk failures (on onl09 (non-critical) RT 2065 and on onlldap (critical) RT 2067)
    • RAID still providing for the file systems in the interim
  • ITD network backups failing (RT 2069), but resolution from ITD may have just come through
  • Machine replacements:
    • emcsc ready (needs either deployed by tomorrow or ITD block needs postponed)
    • emcspin ready (Monday)
    • starutilities awaiting C-AD software

 

FastOffline [Jerome]:

  • Many files coming into the system, but all are tests, pulsers, etc., so not being processed
  • Some missing detector setup, beam energy, etc. information: could mean misconfiguration, or potentially not properly propagated data

 

QA [Gene]:

  • OnlinePlots switched back to newer EVP server
    • Switching steps still undocumented
  • Pre-combining of FastOffline files still needs tested and configured to be the default

 

Run 11 preparation meeting #6

2010-12-17 14:00
2010-12-17 15:00
America/New York
Friday, 17 December 2010
1-189, EVO, at 19:00 (GMT), duration : 01:00

 Minutes

Attendees: Kefeng, Wayne B., Leve H., Gene V.B., Dmitry A.

 

ShiftLog [Leve]:

  • Manual changes have been reviewed (by Jerome) and committed
  • Re-deployed

 

Databases [Dmitry]:

  • cdev enabled (by C-AD), so parameter propagation has begun
    • Some data is still empty (e.g. beam species) and may not be filled for several weeks (when beams actually start)
      • Might be possible to enter some default values for now
    • Numbers are meaningless, so not useful for FastOffline testing
  • Testing new hardware nodes now (2 are available)
    • Configuration of new nodes in discussion

 

Online Infrastructure [Wayne]:

  • STAR login environment needs to be able to handle SL5.5 (which was installed on some nodes)
    • Will follow up with Jerome
  • Online linux pool is losing roughly 1 node a day (rebooting)
    • Hit a few with network scans, but response was OK, so no clues
  • Spin request (for Pibero):
    • User has access to the nodes and is waiting for directory structures and mount points
      • Expect to complete within a few days
  • Condor installation on for rterm-like access to online pool from gateway
    • Installed on the gateway
    • Components for the pool nodes is to be done
  • Slowness on evp
    • Only occurred yesterday (Dec. 16)
    • Nothing obvious, but suspicious of AFS, and coincided with mock data transfer challenge from counting house to HPSS
    • Paths forward discussed:
      • Remove outside AFS dependence using an online repository
      • Remove AFS dependence using local codes on evp
      • Write a new tool better identify AFS issues (i.e. more proactive than 'fs checkservers')
    • Problem doesn't seem to persist, not a priority for now (no action)
  • Yury G. requested networking support for the east FPD/FMS rack
    • Just extends the starp network geographically
    • Needs to be a fiber connection for proper grounding
  • Webserver replacement
    • Two redhat 5 machines given the ITD thumbs up
    • Shared filespaces in progress
    • No services started yet
    • Request for an account to test tomcat (Leve)

 

QA [Gene]:

  • Demo of the Offline QA with reference histogram comparison
    • Missing some features for flexibly defining reference histogram sets (will work on over the holidays)
    • Minor suggestion made for improving the "waiting..." display

Run 11 preparation meeting #7

2011-01-07 14:00
2011-01-07 15:00
America/New York
Friday, 7 January 2011
1-189, EVO, at 19:00 (GMT), duration : 01:00

 Minutes

Attendees: Leve H., Dmitry A., Jerome L., Wayne B., Gene V.B.

 

Databases [Dmitry]:

  • All DB collectors running now
    • Only TPC voltages absent due to no data yet
  • Slow Controls Archiver not yet running to avoid collecting value-less data
    • With Run start imminent, this needs to get going
    •  Dmitry will follow up with Yury G.
  • Potential to benefit from an unused EPICS feature to read multiple channels in one request (we have been reading one channel at a time across the board), which could cut down the overhead in obtaining data
    • Unknown why this wasn't previously used, so testing with caution to learn (perhaps the multiple channel data comes in a burst which could fail for some reason)
  • New shift sign-up release is imminent (possibly today)
    • Only minor features getting final adjustments (features have been [node:19902 "previously presented"])
  • RunLog working fine: new runs appearing, but all being marked bad
    • Jerome notes that bad runs won't come through FastOffline, which has been turned on already
  • New DB nodes
    • Tested and ready for use
    • New configuration (relevant for FastOffline use) not yet in place (Jerome and Dmitry will work this out)
  • Online DB plots working and logs show usage
    • Only unavailable quantities are those collected from the Slow Controls Archiver

 

ShiftLog [Leve]:

  • New manual is printed out and in place at the counting house
  • ShiftLeader desktop computer has all the necessary icons, correctly linked and tested, including making a ShiftLog entry
  • New web server not critical (stable operation on old web server)

 

Online infrastructure [Wayne & Jerome]:

  • OS upgrade
    • One critical: a replacement node for usage in monitoring chilled water (old machine is Windows 2000) is awaiting ITD processing.
      • Follow-up with ITD on Monday if nothing transpires by then
    • Other machines in the queue ("bond", "beatrice", "l3disp")
  • Online web server replacement
    • Currently replicating old machine on new one: file copying and version checking
      • ShiftLog request to continue with same Tomcat version (a known quantity) instead of upgrading in hopes of improved stability
    • Spin resource request involves storage to be delivered on new server
  • Uncertainty in status of UPS services for computers at the experiment
    • Testing today before the Run starts
  • Network request for east FMS/FPD racks completed
  • Online linux pool has been stable for the past ~3 weeks (after a couple weeks of apparently random reboots)

 

FastOffline [Jerome]:

  • Ready, running, and waiting for data

 

QA [Gene]:

  • OnlinePlots still running on older evp machine (but stable)
    • Will look into moving it back over to the new node
    • Need Jeff L. to flip some switches (and document it)
  • Offline QA
    • Finished implementing tools to update just specific histograms in reference set
    • QA Shift set of histograms will be flushed next week and started anew
      • Histograms only go into the set if given a description, and only stay in the set if a reference is set

 

We are expecting some collider operations imminently, so data will likely flow through the entire system over the next week.

 

Software and Computing Phone Meeting

2011-01-05 12:00
2011-01-05 13:00
America/New York
Wednesday, 5 January 2011
1-189, EVO, at 17:00 (GMT), duration : 01:00
TimeTalkPresenter
12:00FMS simulation, open request and readiness ( 00:20 ) 2 filesPibero Djawotho (TAMU)
12:20Production status ( 00:15 ) 0 filesLidia Didenko
12:35AOB ( 00:20 ) 0 filesAll (All)

h+/h- in Run 10 and beyond

2010-12-17 15:00
2010-12-17 16:00
America/New York
Friday, 17 December 2010
EVO, at 20:00 (GMT), duration : 01:00

Software and Computing phone meeting

2010-12-15 12:00
2010-12-15 13:00
America/New York
Wednesday, 15 December 2010
1-189, EVO, at 17:00 (GMT), duration : 01:00
TimeTalkPresenter
12:00Multi-site Data transfer project - status ( 00:20 ) 1 fileMichal Zerola (NPI / ASCR)
12:20Year 2011 geometry status and production intents for Y11 ( 00:15 ) 0 filesJason Webb (BNL)
12:35Production status and stats ( 00:15 ) 0 filesJerome Lauret (BNL)
12:50AOB ( 00:10 ) 0 filesAll (All)

Run 11 preparation meeting #5

2010-12-10 14:00
2010-12-10 15:00
America/New York
Friday, 10 December 2010
1-189, at 19:00 (GMT), duration : 01:00

Attendees: Leve H., Dmitry A., Jérôme L., Wayne B., Gene V.B., Jeff L.

 

Minutes:

Databases [Dmitry]:

  • Conditions collectors started (all except beam information related ones)
    • Everything running appears to be OK and status can be checked from web
    • Beam info needs cdev, provided by C-AD once their operations get going for the Run (also includes magnetic field setting; expect data around the time we turn on the STAR magnet)
      • Absence of this data means RunLog does not finalize runs and information is not migrated to Offline, but this is expected to be OK for now for anyone doing online-only tests (requested to post a note to rts-hn about this)
    • No migration issues expected, but TPC anode voltage recording practices need to be ironed out in the context of the Run 11 data (discuss with Maxim)
  • Online DBs flushed
    • As expected, a few failed links and missing web pages were noted (by users too) and fixed
  • Once cdev is running, will be ready for turn on / testing of FastOffline

 

ShiftLog [Leve]:

  • As expected, excluded from and not affected by DB flush
  • Manual updated for instructions on changing run status (may receive further review & changes from Jerome)

 

Online nodes [Wayne]:

  • AFS has been stable on the newer evp machine since last Friday
    • Jeff's been using the machine (and AFS) without incident
    • Will likely switch OnlinePlots back to this machine (from evp2, the older machine) next week
      • Documentation requested for switching between these machines
  • Replacement of dean at the stage of OS installation on new hardware
    • Timeline for replacement should be well before Run starts, as disk space on the new machine has been requested for online use
  • Trigger commissioning is a high priority part of the request for online node resources (Wayne will contact Pibero directly)
  • Nothing to report on OS upgrades

 

Online QA - jevp demo [Jeff]:

  • Code
    • Subsystems have an xxxBuilder class (e.g. tpcBuilder, eemcBuilder, etc.) which inherits from a jevp base class for plots
    • Plot objects can have multiple histograms added to them, as well as other graphical objects like lines and circles (as long as they are of some "component" type?)
  • Configuration
    • Editor allows hierarchical arrangement of plots into tabs and sub-tabs, under larger Sets (e.g. Shift, 2009, ESMD)
      • Features drag & drop for re-arrangement
    • Tabs can have collective properties for all plots underneath (such as setting all to have the same maximum)
    • Plots can have individual properties (such as log y scale)
  • Running
    • One server starts up several builders as separate processes
    • Data is sent to the builders from a file or an event pool
    • Presenter runs separately to display plots
      • Similar basic look & feel to current OnlinePlots presenter
      • Clicking on a plot brings up a reference side-by-side
        • Drag & drop to add new reference, or select a different reference from panel of choices
        • Suggested: tag references with run from which they're made, and timestamp (already includes a comment)
        • Deleting references is simple, but not so simple as to be done accidentally
        • Mild concerns to be wary of reference bloat and organization as things evolve at the experiment (including collision species/energies change)
    • PDFs generated and uploaded to RunLog DB for each Set
      • Include table of contents
      • Concerns of space this will consume on the onldb server and redundancy of plots
        • Run 10 PDFs used more than half the available disk space, and it is reasonable to expect a doubling (or more) of the required space for Run 11
        • Wayne & Dmitry will assess the current storage on onldb [post-meeting statement that space may be fine; follow up next week]
        • Requirements still need to be more accurately defined
          • Will depend on what gets axed/kept from review by subsystems & trigger board
          • Would be helped by real data, but that's a bit late

Next week: demo of Offline QA for AutoQA with reference histograms [Gene]

Software and Computing phone meeting

2010-12-08 12:00
2010-12-08 13:00
America/New York
Wednesday, 8 December 2010
1-189, EVO, at 17:00 (GMT), duration : 01:00
TimeTalkPresenter
12:00Summary of the h+/h- meeting ( 00:20 ) 0 filesGrant Webb (UKY)
12:20Production status ( 00:15 ) 0 filesLidia Didenko (BNL)
12:35AOB ( 00:20 ) 0 filesAll (All)

Review of Past Issues and Current Understanding


Talk time : 15:20, Duration : 00:20